Find the x and y intercepts and graph the equation plotting the intercepts. Please show all of your work. Submit your graph to the dropbox. 4x-7y=-28
Answer by rfer(16322)

4*0-7y=-28
y=-28/-7
y=4
(0,4)
----------
4x-7*0=-28
4x=-28
x=-28/4
x=-7
(-7,0)